charge q 1 = 3.2 × 10-6 C is placed at the origin and charge q 2 = -5.1 × 10-6 C is placed on the x-axis, at x = -0.20 m. Where along the positive x-axis can a third charge Q = -8.3 µC be placed such that the resultant force on this third charge is zero?
Answer #1

Қx g + х + kg (х+ 0.2) = + 0.2 )2 S. 1xb-6 = x+ O2 = 1.26 x = x = 0.761 M » -